# Policy Contrastive Decoding for Robotic Foundation Models

### v3

@@ 3.2 Policy Contrastive Decoding (PCD)
KDE-based Probabilistic Modeling. We address the second question using KDE-based Probabilistic Modeling (KDE-PM), which approximates action probability distributions for diffusion-based policies through kernel density estimation (KDE) (Wkeglarczyk, 2018). Specifically, we first use the pretrained diffusion-based policy to sample NN candidate action predictions: $\{\bm{a}_{i}(j)\}_{j=1}^{N}$, where $\bm{a}_{i}(j)=[{a}_{1}(j),...,{a}_{M}(j)]\sim\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})$. With the assumption that all action dimensions are independent, i.e., $\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})=\prod\limits_{t=1}^{M}\pi_{\theta}(a_{t}|l,\bm{o}_{i})$, we calculate an approximated probability distribution for each action dimension of the action $\bm{a}_{i}$ via KDE:
@@
where $C^{\prime}$ is a normalization constant and $\mathcal{K}(\cdot)$ indicates a Gaussian kernel $\mathcal{K}(u)=\frac{1}{\sqrt{2\pi}}e^{-\frac{u^{2}}{2}}$, $b$ is the bandwidth parameter controlling the smoothness of the distribution, and $\{a_{t}(j)\}_{j=1}^{N}$ are the NN candidate actions for the $t$-th dimension of the action $\bm{a}_{i}$, generated from the diffusion process. Similarly, we can derive an action probability distribution for $a_{t}$ conditioned on object-masked $\hat{\bm{o}}_{i}$, i.e., $\pi_{\theta}(a_{t}|l,\hat{\bm{o}}_{i})$. With the independence assumption, from $\pi_{\theta}(a_{t}|l,\bm{o}_{i})$ and $\pi_{\theta}(a_{t}|l,\hat{\bm{o}}_{i})$ we can compute the joint action distributions, which are then used to compute the contrastive action probability $\pi_{\theta}^{*}(\bm{a}_{i}|l,\bm{o}_{i})$ using Eq. (2). In this way, our PCD is compatible with diffusion-based robot policies.
@@ 4 Experiments
We answer the first question in Sections 4.1 and 4.2, the second question in Section 4.3, and the third question in Section $4.$ Detailed descriptions of the baseline robot policies and the evaluation tasks used in simulation and real-world environments are presented in Appendix C.
@@ 4.1 Simulation Experiments
Experimental Setup. We conduct experiments on top of three diverse robot policies, including the autoregressive policy OpenVLA (Kim et al., 2024), and the diffusion-based policies Octo (base) (M. et al., 2024) and $\pi_{0}$ (Black et al., 2024). We treat these policies as black-boxes, applying PCD solely based on the output action probability distributions for OpenVLA and the output actions for Octo and $\pi_{0}$. For our PCD, we annotate the target object in the initial observation using three manners: artificial Point and Box prompts, and automatic detection results from Grounding DINO (Liu et al., 2024b). The number of the sampled noise vectors NN in KDE-PM is set to 24 for Octo and $\pi_{0}$. We perform ablation studies on the hyperparameter $\alpha$ in Eq. (2), as well as on the object detection strategies and object inpainting paradigms of the devised Track2Mask module in Section 4.3.

### v4

@@ 3.2 Policy Contrastive Decoding (PCD)
KDE-based Probabilistic Modeling. We address the second question using KDE-based Probabilistic Modeling (KDE-PM), which approximates action probability distributions for diffusion-based policies through kernel density estimation (KDE) (Wkeglarczyk, 2018). Specifically, we first use the pretrained diffusion-based policy to sample $N$ candidate action predictions: $\{\bm{a}_{i}(j)\}_{j=1}^{N}$, where $\bm{a}_{i}(j)=[{a}_{1}(j),...,{a}_{M}(j)]\sim\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})$. With the assumption that all action dimensions are independent, i.e., $\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})=\prod\limits_{t=1}^{M}\pi_{\theta}(a_{t}|l,\bm{o}_{i})$, we calculate an approximated probability distribution for each action dimension of the action $\bm{a}_{i}$ via KDE:
@@
where $C^{\prime}$ is a normalization constant and $\mathcal{K}(\cdot)$ indicates a Gaussian kernel $\mathcal{K}(u)=\frac{1}{\sqrt{2\pi}}e^{-\frac{u^{2}}{2}}$, $b$ is the bandwidth parameter controlling the smoothness of the distribution, and $\{a_{t}(j)\}_{j=1}^{N}$ are the $N$ candidate actions for the $t$-th dimension of the action $\bm{a}_{i}$, generated from the diffusion process. Similarly, we can derive an action probability distribution for $a_{t}$ conditioned on object-masked $\hat{\bm{o}}_{i}$, i.e., $\pi_{\theta}(a_{t}|l,\hat{\bm{o}}_{i})$. With the independence assumption, from $\pi_{\theta}(a_{t}|l,\bm{o}_{i})$ and $\pi_{\theta}(a_{t}|l,\hat{\bm{o}}_{i})$ we can compute the joint action distributions, which are then used to compute the contrastive action probability $\pi_{\theta}^{*}(\bm{a}_{i}|l,\bm{o}_{i})$ using Eq. (2). In this way, our PCD is compatible with diffusion-based robot policies.
@@ 4 Experiments
We answer the first question in Sections 4.1 and 4.2, the second question in Section 4.3, and the third question in Section 4.4. Detailed descriptions of the baseline robot policies and the evaluation tasks used in simulation and real-world environments are presented in Appendix C.
@@ 4.1 Simulation Experiments
Experimental Setup. We conduct experiments on top of three diverse robot policies, including the autoregressive policy OpenVLA (Kim et al., 2024), and the diffusion-based policies Octo (base) (M. et al., 2024) and $\pi_{0}$ (Black et al., 2024). We treat these policies as black-boxes, applying PCD solely based on the output action probability distributions for OpenVLA and the output actions for Octo and $\pi_{0}$. For our PCD, we annotate the target object in the initial observation using three manners: artificial Point and Box prompts, and automatic detection results from Grounding DINO (Liu et al., 2024b). The number of the sampled noise vectors $N$ in KDE-PM is set to 24 for Octo and $\pi_{0}$. We perform ablation studies on the hyperparameter $\alpha$ in Eq. (2), as well as on the object detection strategies and object inpainting paradigms of the devised Track2Mask module in Section 4.3.